OpenStudy (anonymous):

Let logbA=3 logbC=2 logbD=5 then what is the value of logb d^2/ c^3a

OpenStudy (whpalmer4):

remember that \[\log_b x = a\]means that \[b^a = x\]

OpenStudy (shiraz14):

@infamous: You mean: logb (d²/c³a)?

OpenStudy (anonymous):

I got c? am I right?

OpenStudy (whpalmer4):

well, let's see: \[\log_b a = 3\]\[b^3 = a\]\[\log_b c = 2\]\[b^2 = c\]\[\log_b d = 5\]\[b^5 = d\] \[\log_b \frac{d^2}{c^3a} = \log_b \frac{(b^5)^2}{(b^2)^3(b^3)} = \]